the number of complex numbers z which satisfy z2 + 2 modz2= 2 is?

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
can u plz post a pic so the ques is clear.
i guess u r askin
z^2 + 2|z|^2= 2
let z= x+iy then we get
x^2 – y^2 + 2xyi + 2(x^2+y^2)= 2
equating real and imaginary parts,
2xy=0 or xy=0 and x^2 – y^2 + 2(x^2+y^2)= 2.
if x=0, then y= ± sqrt2 
if y=0, then x= ± sqrt(2/3)
so, there are 4 such complex numbers.
kindly approve :)